OpenStudy (anonymous):

FAN AND MEDAL y = 6x − 4 y = 5x − 3 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations.

OpenStudy (dan815):

okay you see y is equal to 2 things with x

OpenStudy (dan815):

if y =6x − 4 and y= 5x − 3 then 6x − 4 = 5x-3 and solve for x

OpenStudy (dan815):

the thinking for this is simple like

OpenStudy (dan815):

y=10+1 y=9+2 10+1 = 9 +2 this ofcourse is true as they are both just saying 11 in different ways

OpenStudy (dan815):

in the same way 6x-4 and 5x-3 are both y just saying it different

OpenStudy (anonymous):

What is the solution to the pair of equations? @dan815

OpenStudy (anonymous):

@DaBest21

OpenStudy (lynfran):

still need help?

OpenStudy (anonymous):

What is the solution to the pair of equations? @LynFran

OpenStudy (lynfran):

are you speaking about this part...Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations.

OpenStudy (anonymous):

no

OpenStudy (lynfran):

ok what are the pair equations ur talking about?

OpenStudy (anonymous):

if its graphed what is the soution as in coordinates

OpenStudy (lynfran):

Did u find what was x already? ...have you graph, the graph already?

OpenStudy (anonymous):

no

OpenStudy (lynfran):

ok we equate the two equation to solve for x i.e 6x-4=5x-3 so, 6x-5x=4-3 and x=1

OpenStudy (lynfran):

so when x=1 to get the y value, we plug x=1 into the original function...i.e eq1 or eq2 to find for y...can you do that?

OpenStudy (anonymous):

can someone just solve this

OpenStudy (mathmate):

@hayespurdy You need to show some efforts on your work! So far you have shown no efforts!

OpenStudy (anonymous):

thats because i have no idea how to solve it @mathmate

OpenStudy (mathmate):

@hayespurdy That's exactly what these nice people are trying to do for you, help you solve \(your\) question. If you have no idea what to do, say exactly which part you do not understand so that others can explain to you. Simply asking for the answer multiple times does not show your willingness to learn. Please participate in the learning process if you want to succeed.

OpenStudy (anonymous):

ok i need you to help me solve for the pair of equations @mathmate

OpenStudy (anonymous):

for the solution

OpenStudy (lynfran):

ok @hayespurdy we had found the value of x earlier just plug in the value of x in one of the given equations and you will get the solution to the problem...

OpenStudy (mathmate):

You are still looking for the solution. I will help you, but tell you ahead of time that I do not give solutions. I leave you to find it.

OpenStudy (anonymous):

ok

OpenStudy (mathmate):

Let's solve, for example, y = 7x − 4 y = 6x − 3 Please do the first step using @dan815 's help, noting that both equations have y on the left-hand side.

OpenStudy (mathmate):

@hayespurdy are you still there?